OpenStudy (anonymous):

What is the potential energy of an electron that is 22.0 cm from a charge of 43.0 nC? How much work is required to move the electron very far from the charge?

OpenStudy (theeric):

I can give you a hint on the first part. Potential energy is (I think opposite) the integral of the force with respect to displacement... Which doesn't mean anything if you're not in calculus. \(U=k\dfrac{q_1\ q_2}r\)

OpenStudy (theeric):

One charge is given, the other charge is that of an electron (which you can probably find in your book).

OpenStudy (theeric):

Good luck! I have to get going!

OpenStudy (anonymous):

yea thats exactly what i did but it wasnt the right answer

OpenStudy (anonymous):

Did @theEric's help make sense to you? I think it was complete.

OpenStudy (anonymous):

the two parts of the question ask the same thing...

OpenStudy (anonymous):

i didnt get the right answer i got -2.81x10^-16 J

OpenStudy (anonymous):

i also tried +2.81x10^-16 J

OpenStudy (anonymous):

It should be. because you need extra energy to move the electron more away from the other. it is the external work.

OpenStudy (anonymous):

so how can i calculate it ? Because this one wouldn't work

OpenStudy (anonymous):

I am doing some quantum problems and dont have time to write equations. I would like @theEric to help you further!

OpenStudy (theeric):

The charge of the electron is about \(-1.602\times10^{-19}\ \rm C\). The other charge is \(43.0\ \rm{nC}=43.0\times10^{-6}\ \rm C\) So \(q_1\ q_2=(-1.602)(10^{-19})(43.0)(10^{-6})=-68.9\times10^{-25}\). Is that what you got to that point?

OpenStudy (theeric):

\(r=22.0\ \rm{cm}=22.0\times10^{-2}\ \rm m\) And there's that. The reason I'm going through this now is because I see all the different units and wonder if that is a point of possible error.

OpenStudy (anonymous):

1n=10^-9m

OpenStudy (theeric):

Coulomb's constant is about \(\large8.988\times10^9\ \rm\frac{N\bullet m^2}{C^2}\) (the dot is just multiplication).

OpenStudy (theeric):

Thank you @Saeeddiscover !!! :) Bad mistake...

OpenStudy (theeric):

The charge of the electron is about \(−1.602×10^{−19}\ \rm C.\) The other charge is \(43.0\ \rm{nC}=43.0\times10^{-\color{red}9}\ \rm C\) So \(q_1\ q_2=(-1.602)(10^{-19})(43.0)(10^{-\color{red}9})(\rm{C^2})=-68.9\times10^{-\color{red}{28}}\ \rm{C^2}\)

OpenStudy (theeric):

Again, \(r=22.0\ \rm{cm}=22.0\times10^{-2}\ \rm m\) and \(k=\large8.988\times10^9\ \rm\frac{N\bullet m^2}{C^2}\)

OpenStudy (theeric):

\[U=k\frac{q_1\ q_2}r\]\(=-28.1\times10^{(9)+(-28)-(-2)}\ \rm J=-28.1\times10^{-17}\ \rm J\\=-2.81\times10^{-16}\ \rm J\) Interesting... And we would say that its potential is positive.... So \(-2.81\times10^{-16}\ \rm J\).

OpenStudy (theeric):

I meant, so \(2.81\times10^{-16}\ \rm J\).

OpenStudy (theeric):

Moving the electron to the positive charge.. The electron will do that naturally, so the electrostatic force does the work. The work done by it is positive. The work done on it is negative. Other than that, I forget the formula. I could derive it, but I don't know if I should trust myself. Logically, it's the opposite the work to move it from where it is (22.0 cm away) to infinity (where the electrostatic force is theoretically \(0\)).

OpenStudy (theeric):

I have to go! Take care!

OpenStudy (anonymous):

ohh i did opposite the Work is negative alright thanks

OpenStudy (theeric):

Di you get them right?

OpenStudy (theeric):

did*

OpenStudy (anonymous):

yea i did!! thanks

OpenStudy (theeric):

Congrats!